∫(1/x + 3/x2 - 4/x3 ) dx



1
Expert's answer
2021-05-27T11:32:13-0400

"\Rightarrow \int(\dfrac{1}{x} + \dfrac{3}{x^2} -\dfrac{4}{x^3})dx"


"\Rightarrow \int( x^{-1} +3 x^{-2}-4x^{-3})dx"


"\Rightarrow ln | x|+3 \dfrac{x^{-2+1}}{(-2+1)} -4 \dfrac{x^{-3+1}}{(-3+1)}"


"\Rightarrow ln |x| -\dfrac{3}{x} + \dfrac{2}{x^2}"